• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name S. Bailey

    Log entry time 23:51:33 on July26,2005

    Entry number 149600

    keyword=Swing Shift Summary

    Swing shift summary:
    Shift workers:
    S. Bailey, K. Grimm, J. Katich
    
    16:00 Still in controlled access.
    16:00 Pedestal Run 3845 (no beam)
    17:00 Moved from 4He target to empty target
    17:08 Pedestal Run 3846, 5uA at the very end
    17:20 Power Permit
    17:30 5 uA, 14x14 raster, spott++
    17:35 Moved from empty target to 4He
    17:45 Increased current to 40 uA, in increments of 10 uA
    17:50 Called MCC to decrease y-positions of both BPMA and BPMB
    17:54 Production Run 3837 
    18:05 Start Compton Run 
    18:40 Septum quenched at 7.7 K (Run 3837), beam off
    19:04 Pedestal Run 3848
    --------------
    While we wait for the septum temperature to drop below 7.8K, MCC will install 
    new optics. The following runs are taken with the new optics.
    19:25 Moved from 4He target to empty target
    19:30 Dithering Test, Run 3849, 5 uA
    19:40 Dithering Test, Run 3850, 30 uA
          Results show no improvement.
    19:45 Called MCC and asked them to back out of the change. No beam.         
    Continue to cool septum temperature.
    19:55 Moved from empty target to 4He target
    20:00 Septum temperature @ 7.8 K. Reset right septum interlock, set         
    right current septum
    20:04 Pedestal Run 3851 (no beam)
    ----------------
    20:45 The right septum temperature is 6.658 K and the right septum          
    current is 278.22 A. We are ready to take production runs.
    20:45 MCC lost the electron gun due to a radiation alarm. They are          
    troubleshooting the problem. 
    20:54 Pedestal Run 3852 (no beam)
    21:05 Beam is back, 5 uA, spot++
    21:15 Increased to 38 uA, in increments of 10 uA
          Interrupted. Beam down to tune Hall C.
    21:20 Beam returns. Increased to 38 uA, in increments of 10 uA.
    21:22 Production Run 3853
    21:22 Start Compton Run 11561
    21:35 The right septum temperature was fastly approaching 7.6K so I         
    dropped beam current from 38 uA to 35 uA.
    22:00 Dropped beam current from 35 uA to 34 uA. 
    22:05 Stopped Compton Run 11561 due to Roc5 problems. Started Compton       Run 
    11562.
    22:13 Production Run 3854
    22:20 Tried to adjust raster size to yield 4 x 4 (suggested by Rob F.)
          I did numerous iterations in both x and y and there was no            
    clear correlation between set points and spot++ points. Back to       14 x 14 
    raster size.
    22:26 Production Run 3855
    22:38 Beam down, box supply needs to be fixed.
    22:40 No response from Happex DAQ. Rebooted DAQ twice. Adaql1               
    rebooted itself. According to Bryan, right spectrometer needs         to 
    reboot. He said to request a controlled access. Rob is             coming in 
    and will do this.
    22:45 Rob enters Hall in Controlled Access